# Safe But Not Sorry: Reducing Over-Conservatism in Safety Critics via Uncertainty-Aware Modulation

### v1

@@ 1. Introduction
Reinforcement Learning (RL) empowers agents to autonomously devise sequential decision-making strategies by interacting with their environment sutton2018reinforcement (36). Its successes span a broad spectrum of domains, ranging from mastering strategic games open-ai-five (8) to enabling sophisticated robotic control loc-behaviours (17, 29), exhibiting capabilities on par with human cognition. Although exploration is the foundation through which an agent learns, it can become particularly problematic in safety-critical domains, such as healthcare and robotics, where physical or ethical risks may occur from unsafe actions. During exploration in these domains, the agent can naturally encounter unfamiliar and potentially hazardous states in which erroneous actions can yield unacceptable consequences.
@@
Addressing this challenge mandates the design of RL agents that can both avoid unsafe behaviours and retain sufficient exploratory capacity to uncover optimal policies concrete-problems (5). Ensuring safety is essential not only during training, when the agent explores and adapts, but also throughout deployment, where its decisions are observable in the real world and directly affect its stakeholders.
@@
Prior safe RL approaches can be classified into external knowledge-based and cost-based. External knowledge approaches leverage formal specifications gross2022cool (15), expert demonstrations li2024guided (22), or interactive oversight saunders2017trial (32) to constrain exploration and prevent unsafe behavior odriozola2023shielded (24). Despite their effectiveness, they often require substantial domain expertise or supervision bastani2018verifiable (6). Cost-based methods, instead, embed safety directly into the learning problem by formulating RL as a constrained optimisation task within the Constrained MDP (CMDP) framework altman1998constrained (4). Within this scheme Lagrangian relaxation stooke2020responsive (35), primal–dual optimisation paternain2022safe (25), and refined objectives satija2020constrained (31, 43) balance reward against cumulative cost. These methods, however, rely on estimates of expected future cost produced by safety critics, whose estimation capability is severely affected by noisy or sparse cost signals srinivasan2020learning (34). In contrast, conservative variants assume that uncertain or underexplored states might incur higher risk than in reality and mitigate underestimation by biasing toward higher costs bharadhwaj2020conservative (12).
@@
Safety critics bharadhwaj2020conservative (12, 34), while central to enforcing safety constraints, often produce distorted gradient signals that hinder learning castellano2023learning (14). By uniformly inflating or flattening cost estimates across large regions of the state–action space, these critics blur the boundary between genuinely unsafe and merely uncertain actions. Consequently, the gradients passed to the policy become weak, saturated, or misaligned with the underlying risk landscape. In already safe regions, inflated cost signals can suppress reward gradients, preventing policy refinement and improvement. Conversely, when the dual variable in the Lagrangian framework is high, the critic’s outputs may saturate, marking nearly all actions as unsafe and collapsing the gradient structure castellano2023learning (14). This collapse leaves the policy without informative guidance, stalling progress and undermining the delicate trade-offs between reward maximisation (performance) and constraint satisfaction (safety).
@@
We address this significant limitation by introducing Uncertain Safety Critic (USC), a novel approach that reduces over-conservatism in safety critics by explicitly considering parameter-space uncertainty in cost estimation. Unlike standard safety critics srinivasan2020learning (34, 12), which often flatten or inflate risk landscapes indiscriminately, USC selectively modulates conservatism proportionally to the critic’s epistemic uncertainty, preserving informative gradients while still erring on the side of caution. Drawing inspiration from recent advances in influence-based uncertainty quantification koh2017understanding (20), USC shapes its conservative bias through uncertainty-weighted objectives and further sharpens predictions via an uncertainty refinement procedure that interpolates poorly represented regions of the state–action space. By doing so, USC produces sharper and more reliable cost maps, enabling policies to achieve scalable and precise reward–cost trade-offs, with gradient errors reduced by $\approx 83\%$ compared to the state-of-the-art, a result that we further substantiate through our theoretical analysis. Moreover, USC reduces safety violations by up to $40\%$ while simultaneously achieving higher rewards that scale effectively with task complexity across nearly all evaluated settings. To the best of our knowledge, USC is the first safety critic method that couples conservative estimation with uncertainty-aware modulation, directly addressing the gradient collapse that undermines existing RL safety critics.
@@ 2. Related Work
Approaches Based on External Knowledge. Several strategies for enabling safe exploration in reinforcement learning leverage prior knowledge or explicit safety constraints gu2024review (16). By extracting the agent’s policy during learning and utilising formal verification techniques, these policies can be verified against predefined safety specifications, providing guarantees on their behaviour before deployment bastani2018verifiable (6, 15). An analogous way to utilise prior knowledge for safe exploration involves using expert demonstrations to initialise the agent’s understanding of task dynamics and safety-critical behaviour li2024guided (22, 28). Similarly, interactive oversight during training allows experts to intervene when the agent is likely to take a dangerous action xu2022look (42, 32, 27), or to design a curriculum that gradually exposes the agent to more challenging or risky scenarios turchetta2020safe (39). An alternative approach employs safety shields, which intercept or correct unsafe actions during execution konighofer2020shield (21, 24). While some shields rely on predefined safety constraints through LTL alshiekh2018safe (2) or probabilistic specifications jansen2020safe (18), this requires domain knowledge; others can be learned online shperberg2022learning (33), partially relaxing the need for prior knowledge. While these techniques can be highly effective, they rely to varying degree on often unavailable or costly expert input, domain knowledge, or structured priors. This motivates an alternative direction: enabling safety via environmental feedback alone, without assuming external supervision or formal task knowledge bethell2024safe (11).
@@
Cost-based Learning Approaches. In this setting, safety is no longer enforced through external supervision or logical constraints, but instead learned through interaction with the environment wachi2024survey (40). By incorporating cost signals into the MDP framework to provide the agent with feedback on risky actions or outcomes (altman2021constrained, 3, p. 21), similar to task feedback via reward, the agent can be trained to balance task performance against cumulative risk, without requiring prior knowledge of what constitutes safe behaviour. A common approach is to formulate the agent’s objective as a constrained optimisation problem, where the goal is to maximise expected reward subject to an upper bound on expected cumulative cost. Dual-based methods that use a Lagrangian relaxation of the constraint are among the most widely adopted approaches in this setting. These methods introduce a cost multiplier to penalise constraint violations during policy learning, effectively transforming the constrained problem into unconstrained by augmenting the reward signal altman1998constrained (4, 35). However, this formulation can lead to instability or poor convergence if the multiplier is not well-tuned or if cost gradients are unreliable. Primal-dual optimisation paternain2022safe (25) jointly updates the policy and dual variables using structured learning rules to guarantee convergence. RCPO tessler2018reward (37) decouples the reward and cost learning signals to avoid interference, using a soft constraint formulation that adjusts the constraint budget based on policy performance. More recently, methods such as Convergent Policy Optimization yu2019convergent (43) and Backwards Value Functions satija2020constrained (31) have introduced reformulations of the CMDP objective that ensure convergence to feasible policies, even under complex constraint landscapes or sparse cost signals. Across all these methods, however, a shared and often under-examined challenge remains: accurately estimating expected cumulative cost. Since policies are trained to balance reward against predicted cost, the reliability and structure of the cost signal, typically learned through a safety critic, play a critical role in both constraint satisfaction and overall performance.
@@
Safety Critic (SC) Approaches. Central to the success of cost-based approaches is the accurate estimation of expected cumulative cost, which informs both the policy’s constraint satisfaction and its trade-off with reward gu2024review (16). This estimation is typically carried out by a safety critic, a value function trained to predict future cost under the current policy srinivasan2020learning (34). In practice, standard critics often struggle with sparse or high-variance cost signals, leading to unstable or inaccurate learning. To mitigate this, recent research has proposed conservative safety critics (CSC) bharadhwaj2020conservative (12) that deliberately overestimate cost across the state-action space. Rather than aiming to estimate the true expected cost, CSC encourages the critic to assign higher values by penalising low-cost predictions for new policy actions. While this approach can reduce the risk of underestimating safety violations, it often produces diffused or inflated cost landscapes that obscure genuine risk boundaries. Consequently, policies trained under CSC may receive weak or saturated gradients, limiting their ability to make fine-grained reward–cost trade-offs. For example, Figure 1 illustrates the predictive cost maps generated by the different safety critic variants in the CarGoal2 task of the Safety Gymnasium test-suite ji2023safety (19). While the SC underestimates hazard boundaries, the CSC agent inflates cost uniformly across the environment space, producing diffuse cost maps that obscure genuine risk regions.
@@ 3. Preliminaries
Markov Decision Process. A Markov Decision Process (MDP) bellman1957markovian (7) is a discrete-time stochastic control framework for sequential decision-making. An MDP is defined as the tuple $M=(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ is the action space, $P(s_{t+1}|s_{t},a_{t})$ denotes the probability of transitioning to state $s_{t+1}$ from state $s_{t}$ after taking action $a_{t}$, $R(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. A policy $\pi:S\to\Delta(A)$ maps states to distributions over actions. Classical solution methods like value and policy iteration bertsekas2008introduction (10) assume full knowledge of the MDP dynamics. MDPs include accepting states, which represent task completion, and non-accepting terminal states, which often signal safety violations. In this work, we consider MDPs with continuous state and action spaces, i.e., $S\subseteq\mathbb{R}^{n}$, $A\subseteq\mathbb{R}^{m}$.
@@
A Constrained Markov Decision Process (CMDP) altman2021constrained (3) extends the standard MDP framework to incorporate safety or resource constraints. Formally, a CMDP is defined as a tuple $M^{\prime}=$($S$, AA, $P$, $R$, $C$, $\chi$, $\gamma$) where ($S$, AA, $P$, $R$, $\gamma$) defines the underlying MDP, $C:S\times A\to\mathbb{R}_{\geq 0}$ is a cost function that assigns a non-negative cost to each state-action pair, and $\chi\in\mathbb{R}_{\geq 0}$ denotes the maximum allowable cost per episode. The objective is to find a policy $\pi^{*}$ that maximises the expected return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ while satisfying the constraint $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}C(s_{t},a_{t})\right]\leq\chi$. CMDPs provide a principled approach to encode safety or performance limits directly into the learning process and are commonly addressed using Lagrangian-based methods or constrained optimisation techniques. In this work, CMDPs are used as a foundation for modelling safe decision-making under soft safety constraints, which are enforced in expectation and become hard when the safety budget $\chi=0$.
@@
Reinforcement Learning. Reinforcement Learning (RL) addresses decision-making in unknown environments modelled as MDPs. The agent seeks an optimal policy $\pi^{*}$ that maximises the expected discounted return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ sutton2018reinforcement (36). The value function $V_{\pi}(s)$ estimates the expected return from state $s$ under policy $\pi$. Regret berry1985bandit (9) quantifies the performance gap between the learned and optimal policies, with typical rates $O(1/\sqrt{E_{\mathit{max}}})$, where $E_{\mathit{max}}$ is the training horizon. Canonical algorithms include Q-learning qlearning (41) and SARSA sarsa (30). Deep Reinforcement Learning (DRL) employs neural networks to approximate policies or value functions in high-dimensional settings. Actor-critic methods sutton2018reinforcement (36) decouple the policy (actor) and value function (critic), enabling concurrent updates. The Deep Deterministic Policy Gradient (DDPG) algorithm ddpg-paper (23) exemplifies this approach in continuous action domains.
@@ 4.1. Uncertainty-Driven Training
USC departs from widely-used failure-based safety formulations that rely on binary constraint signals $\{0,1\}$ bharadhwaj2020conservative (12, 34). Instead, USC operates under a continuous cost-based setting achiam2017constrained (1), where each transition incurs a real-valued cost signal $C(s,a)\in\mathbb{R}_{+}$. This formulation enables the safety critic to model both hard and soft constraint violations and is particularly well-suited to environments where constraint violations are graded, such as collision intensity, distance to hazard, or control effort. The safety critic predicts the expected cumulative cost under the current policy $\pi$. Unlike binary failure critics, the continuous cost formulation provides denser supervision, captures meaningful risk gradients essential for scalable safety under tighter constraints, and avoids ambiguity in defining failure, enhancing applicability across diverse safety settings.
The safety critic $Q_{C}$ estimates the cost of a particular action in a given state, and these estimates can be used to guide safer exploration. An ideal safety critic should yield accurate and conservative estimates as well as sharp and informative gradients. To create such a critic, we incorporate upper bound estimates on $Q_{C}$ to overestimate alternative actions from the policy and weight them based on their confidence using Gauss-Newton influence koh2017understanding (20), thus preventing overestimation of well-predicted data. We adopt Gauss-Newton over alternative influence functions because it provides stable, interpretable, and scalable uncertainty estimates using only first-order gradients and a positive semi-definite curvature botev2017practical (13).
@@
The log-exp-sum (LSE) penalty, used in Equation 3 and standard in CSC bharadhwaj2020conservative (12), is more effective at encouraging conservatism when the $m$ alternative actions are sampled uniformly over the action space, rather than from the current policy bharadhwaj2020conservative (12). If the policy aims to avoid high-cost regions, then the uniform sampling is more likely to draw from high-cost out-of-distribution actions. Formally, for any fixed state $s$, the conservative term is strictly larger under uniform scaling.
@@
where $\eta$ is the learning rate for the safety critic. In parallel, the reward critic $Q_{R}$ is trained and updated as standard ddpg-paper (23). The actor is optimised with a Lagrangian objective that balances reward maximisation, cost penalisation, and stability under policy shifts. Its loss is given by
@@
All updates described above are embedded within the standard RL loop (Algorithm 1 with the USC extension). At each step, the actor proposes a candidate action, which is then screened by the safety critic. If the predicted cost exceeds that safety tolerance $\varepsilon_{\text{safe}}$, additional candidate actions are produced using Ornstein–Uhlenbeck noise ddpg-paper (23). Among these candidates, the action with the lowest predicted cost is selected, provided it satisfies the tolerance $\varepsilon_{\text{safe}}$; otherwise, the least risky action is chosen. This mechanism ensures that the policy not only maximises reward but also respects the safety critic’s guidance. Transitions are stored in the replay buffer, from which the actor and both critics are iteratively updated. The dual variable $\lambda$ is dynamically updated to maintain the cost budget. In this way, the safety critic is integral to both training and execution, shaping policy improvement while directly filtering unsafe actions.

### v2

@@ 1. Introduction
Reinforcement Learning (RL) empowers agents to autonomously devise sequential decision-making strategies by interacting with their environment sutton2018reinforcement. Its successes span a broad spectrum of domains, ranging from mastering strategic games open-ai-five to enabling sophisticated robotic control loc-behaviours; massively-drl, exhibiting capabilities on par with human cognition. Although exploration is the foundation through which an agent learns, it can become particularly problematic in safety-critical domains, such as healthcare and robotics, where physical or ethical risks may occur from unsafe actions. During exploration in these domains, the agent can naturally encounter unfamiliar and potentially hazardous states in which erroneous actions can yield unacceptable consequences.
@@
Addressing this challenge mandates the design of RL agents that can both avoid unsafe behaviours and retain sufficient exploratory capacity to uncover optimal policies concrete-problems. Ensuring safety is essential not only during training, when the agent explores and adapts, but also throughout deployment, where its decisions are observable in the real world and directly affect its stakeholders.
@@
Prior safe RL approaches can be classified into external knowledge-based and cost-based. External knowledge approaches leverage formal specifications gross2022cool, expert demonstrations li2024guided, or interactive oversight saunders2017trial to constrain exploration and prevent unsafe behavior odriozola2023shielded. Despite their effectiveness, they often require substantial domain expertise or supervision bastani2018verifiable. Cost-based methods, instead, embed safety directly into the learning problem by formulating RL as a constrained optimisation task within the Constrained MDP (CMDP) framework altman1998constrained. Within this scheme Lagrangian relaxation stooke2020responsive, primal–dual optimisation paternain2022safe, and refined objectives satija2020constrained; yu2019convergent balance reward against cumulative cost. These methods, however, rely on estimates of expected future cost produced by safety critics, whose estimation capability is severely affected by noisy or sparse cost signals srinivasan2020learning. In contrast, conservative variants assume that uncertain or underexplored states might incur higher risk than in reality and mitigate underestimation by biasing toward higher costs bharadhwaj2020conservative.
@@
Safety critics bharadhwaj2020conservative; srinivasan2020learning, while central to enforcing safety constraints, often produce distorted gradient signals that hinder learning castellano2023learning. By uniformly inflating or flattening cost estimates across large regions of the state–action space, these critics blur the boundary between genuinely unsafe and merely uncertain actions. Consequently, the gradients passed to the policy become weak, saturated, or misaligned with the underlying risk landscape. In already safe regions, inflated cost signals can suppress reward gradients, preventing policy refinement and improvement. Conversely, when the dual variable in the Lagrangian framework is high, the critic’s outputs may saturate, marking nearly all actions as unsafe and collapsing the gradient structure castellano2023learning. This collapse leaves the policy without informative guidance, stalling progress and undermining the delicate trade-offs between reward maximisation (performance) and constraint satisfaction (safety).
@@
We address this significant limitation by introducing Uncertain Safety Critic (USC), a novel approach that reduces over-conservatism in safety critics by explicitly considering parameter-space uncertainty in cost estimation. Unlike standard safety critics srinivasan2020learning; bharadhwaj2020conservative, which often flatten or inflate risk landscapes indiscriminately, USC selectively modulates conservatism proportionally to the critic’s epistemic uncertainty, preserving informative gradients while still erring on the side of caution. Drawing inspiration from recent advances in influence-based uncertainty quantification koh2017understanding, USC shapes its conservative bias through uncertainty-weighted objectives and further sharpens predictions via an uncertainty refinement procedure that interpolates poorly represented regions of the state–action space. By doing so, USC produces sharper and more reliable cost maps, enabling policies to achieve scalable and precise reward–cost trade-offs, with gradient errors reduced by $\approx 83\%$ compared to the state-of-the-art, a result that we further substantiate through our theoretical analysis. Moreover, USC reduces safety violations by up to $40\%$ while simultaneously achieving higher rewards that scale effectively with task complexity across nearly all evaluated settings. To the best of our knowledge, USC is the first safety critic method that couples conservative estimation with uncertainty-aware modulation, directly addressing the gradient collapse that undermines existing RL safety critics.
@@ 2. Related Work
Approaches Based on External Knowledge. Several strategies for enabling safe exploration in reinforcement learning leverage prior knowledge or explicit safety constraints gu2024review. By extracting the agent’s policy during learning and utilising formal verification techniques, these policies can be verified against predefined safety specifications, providing guarantees on their behaviour before deployment bastani2018verifiable; gross2022cool. An analogous way to utilise prior knowledge for safe exploration involves using expert demonstrations to initialise the agent’s understanding of task dynamics and safety-critical behaviour li2024guided; ramirez2022model. Similarly, interactive oversight during training allows experts to intervene when the agent is likely to take a dangerous action xu2022look; saunders2017trial; plaut2025asking, or to design a curriculum that gradually exposes the agent to more challenging or risky scenarios turchetta2020safe. An alternative approach employs safety shields, which intercept or correct unsafe actions during execution konighofer2020shield; odriozola2023shielded. While some shields rely on predefined safety constraints through LTL alshiekh2018safe or probabilistic specifications jansen2020safe, this requires domain knowledge; others can be learned online shperberg2022learning, partially relaxing the need for prior knowledge. While these techniques can be highly effective, they rely to varying degree on often unavailable or costly expert input, domain knowledge, or structured priors. This motivates an alternative direction: enabling safety via environmental feedback alone, without assuming external supervision or formal task knowledge bethell2024safe.
@@
Cost-based Learning Approaches. In this setting, safety is no longer enforced through external supervision or logical constraints, but instead learned through interaction with the environment wachi2024survey. By incorporating cost signals into the MDP framework to provide the agent with feedback on risky actions or outcomes (altman2021constrained, p. 21), similar to task feedback via reward, the agent can be trained to balance task performance against cumulative risk, without requiring prior knowledge of what constitutes safe behaviour. A common approach is to formulate the agent’s objective as a constrained optimisation problem, where the goal is to maximise expected reward subject to an upper bound on expected cumulative cost. Dual-based methods that use a Lagrangian relaxation of the constraint are among the most widely adopted approaches in this setting. These methods introduce a cost multiplier to penalise constraint violations during policy learning, effectively transforming the constrained problem into unconstrained by augmenting the reward signal altman1998constrained; stooke2020responsive. However, this formulation can lead to instability or poor convergence if the multiplier is not well-tuned or if cost gradients are unreliable. Primal-dual optimisation paternain2022safe jointly updates the policy and dual variables using structured learning rules to guarantee convergence. RCPO tessler2018reward decouples the reward and cost learning signals to avoid interference, using a soft constraint formulation that adjusts the constraint budget based on policy performance. More recently, methods such as Convergent Policy Optimization yu2019convergent and Backwards Value Functions satija2020constrained have introduced reformulations of the CMDP objective that ensure convergence to feasible policies, even under complex constraint landscapes or sparse cost signals. Across all these methods, however, a shared and often under-examined challenge remains: accurately estimating expected cumulative cost. Since policies are trained to balance reward against predicted cost, the reliability and structure of the cost signal, typically learned through a safety critic, play a critical role in both constraint satisfaction and overall performance.
@@
Safety Critic (SC) Approaches. Central to the success of cost-based approaches is the accurate estimation of expected cumulative cost, which informs both the policy’s constraint satisfaction and its trade-off with reward gu2024review. This estimation is typically carried out by a safety critic, a value function trained to predict future cost under the current policy srinivasan2020learning. In practice, standard critics often struggle with sparse or high-variance cost signals, leading to unstable or inaccurate learning. To mitigate this, recent research has proposed conservative safety critics (CSC) bharadhwaj2020conservative that deliberately overestimate cost across the state-action space. Rather than aiming to estimate the true expected cost, CSC encourages the critic to assign higher values by penalising low-cost predictions for new policy actions. While this approach can reduce the risk of underestimating safety violations, it often produces diffused or inflated cost landscapes that obscure genuine risk boundaries. Consequently, policies trained under CSC may receive weak or saturated gradients, limiting their ability to make fine-grained reward–cost trade-offs. For example, Figure 1 illustrates the predictive cost maps generated by the different safety critic variants in the CarGoal2 task of the Safety Gymnasium test-suite ji2023safety. While the SC underestimates hazard boundaries, the CSC agent inflates cost uniformly across the environment space, producing diffuse cost maps that obscure genuine risk regions.
@@ 3. Preliminaries
Markov Decision Process. A Markov Decision Process (MDP) bellman1957markovian is a discrete-time stochastic control framework for sequential decision-making. An MDP is defined as the tuple $M=(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ is the action space, $P(s_{t+1}|s_{t},a_{t})$ denotes the probability of transitioning to state $s_{t+1}$ from state $s_{t}$ after taking action $a_{t}$, $R(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. A policy $\pi:S\to\Delta(A)$ maps states to distributions over actions. Classical solution methods like value and policy iteration bertsekas2008introduction assume full knowledge of the MDP dynamics. MDPs include accepting states, which represent task completion, and non-accepting terminal states, which often signal safety violations. In this work, we consider MDPs with continuous state and action spaces, i.e., $S\subseteq\mathbb{R}^{n}$, $A\subseteq\mathbb{R}^{m}$.
@@
A Constrained Markov Decision Process (CMDP) altman2021constrained extends the standard MDP framework to incorporate safety or resource constraints. Formally, a CMDP is defined as a tuple $M^{\prime}=$($S$, AA, $P$, $R$, $C$, $\chi$, $\gamma$) where ($S$, AA, $P$, $R$, $\gamma$) defines the underlying MDP, $C:S\times A\to\mathbb{R}_{\geq 0}$ is a cost function that assigns a non-negative cost to each state-action pair, and $\chi\in\mathbb{R}_{\geq 0}$ denotes the maximum allowable cost per episode. The objective is to find a policy $\pi^{*}$ that maximises the expected return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ while satisfying the constraint $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}C(s_{t},a_{t})\right]\leq\chi$. CMDPs provide a principled approach to encode safety or performance limits directly into the learning process and are commonly addressed using Lagrangian-based methods or constrained optimisation techniques. In this work, CMDPs are used as a foundation for modelling safe decision-making under soft safety constraints, which are enforced in expectation and become hard when the safety budget $\chi=0$.
@@
Reinforcement Learning. Reinforcement Learning (RL) addresses decision-making in unknown environments modelled as MDPs. The agent seeks an optimal policy $\pi^{*}$ that maximises the expected discounted return $\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ sutton2018reinforcement. The value function $V_{\pi}(s)$ estimates the expected return from state $s$ under policy $\pi$. Regret berry1985bandit quantifies the performance gap between the learned and optimal policies, with typical rates $O(1/\sqrt{E_{\mathit{max}}})$, where $E_{\mathit{max}}$ is the training horizon. Canonical algorithms include Q-learning qlearning and SARSA sarsa. Deep Reinforcement Learning (DRL) employs neural networks to approximate policies or value functions in high-dimensional settings. Actor-critic methods sutton2018reinforcement decouple the policy (actor) and value function (critic), enabling concurrent updates. The Deep Deterministic Policy Gradient (DDPG) algorithm ddpg-paper exemplifies this approach in continuous action domains.
@@ 4.1. Uncertainty-Driven Training
USC departs from widely-used failure-based safety formulations that rely on binary constraint signals $\{0,1\}$ bharadhwaj2020conservative; srinivasan2020learning. Instead, USC operates under a continuous cost-based setting achiam2017constrained, where each transition incurs a real-valued cost signal $C(s,a)\in\mathbb{R}_{+}$. This formulation enables the safety critic to model both hard and soft constraint violations and is particularly well-suited to environments where constraint violations are graded, such as collision intensity, distance to hazard, or control effort. The safety critic predicts the expected cumulative cost under the current policy $\pi$. Unlike binary failure critics, the continuous cost formulation provides denser supervision, captures meaningful risk gradients essential for scalable safety under tighter constraints, and avoids ambiguity in defining failure, enhancing applicability across diverse safety settings.
The safety critic $Q_{C}$ estimates the cost of a particular action in a given state, and these estimates can be used to guide safer exploration. An ideal safety critic should yield accurate and conservative estimates as well as sharp and informative gradients. To create such a critic, we incorporate upper bound estimates on $Q_{C}$ to overestimate alternative actions from the policy and weight them based on their confidence using Gauss-Newton influence koh2017understanding, thus preventing overestimation of well-predicted data. We adopt Gauss-Newton over alternative influence functions because it provides stable, interpretable, and scalable uncertainty estimates using only first-order gradients and a positive semi-definite curvature botev2017practical.
@@
The log-exp-sum (LSE) penalty, used in Equation 3 and standard in CSC bharadhwaj2020conservative, is more effective at encouraging conservatism when the $m$ alternative actions are sampled uniformly over the action space, rather than from the current policy bharadhwaj2020conservative. If the policy aims to avoid high-cost regions, then the uniform sampling is more likely to draw from high-cost out-of-distribution actions. Formally, for any fixed state $s$, the conservative term is strictly larger under uniform scaling.
@@
where $\eta$ is the learning rate for the safety critic. In parallel, the reward critic $Q_{R}$ is trained and updated as standard ddpg-paper. The actor is optimised with a Lagrangian objective that balances reward maximisation, cost penalisation, and stability under policy shifts. Its loss is given by
@@
All updates described above are embedded within the standard RL loop (Algorithm 1 with the USC extension). At each step, the actor proposes a candidate action, which is then screened by the safety critic. If the predicted cost exceeds that safety tolerance $\varepsilon_{\text{safe}}$, additional candidate actions are produced using Ornstein–Uhlenbeck noise ddpg-paper. Among these candidates, the action with the lowest predicted cost is selected, provided it satisfies the tolerance $\varepsilon_{\text{safe}}$; otherwise, the least risky action is chosen. This mechanism ensures that the policy not only maximises reward but also respects the safety critic’s guidance. Transitions are stored in the replay buffer, from which the actor and both critics are iteratively updated. The dual variable $\lambda$ is dynamically updated to maintain the cost budget. In this way, the safety critic is integral to both training and execution, shaping policy improvement while directly filtering unsafe actions.